Home‎ > ‎Crew 133‎ > ‎

109-CommandersReport

Commander Report
01/09/2014
Paula Crock
 
The crew had a quiet in-Hab day today to work on project and do some cleaning.  The lab floor is
cleaner than I’ve seen it before; too bad we’ll likely get it tracked up tomorrow after getting back
from the EVA.  Komorowski took two more of the crew through his telemedicine training, Jessup worked
on the cabling for the radio telescope, Diaz-Rubin and Gartrelle went through IMO charts to
determine which to use for meteor observations while I repaired one broken helmet temperature logger
and readied another one for the EVA tomorrow.  Starting tomorrow we should begin to see data coming
in on some of our projects so expect science reports heading Earthside soon.
 
With our water supply system offline until Tuesday, the crew’s discussion this evening is how they will get
through the next four days with the water currently in storage.  I know crews are quite capable of
managing on this volume without much sacrifice, just extra diligence, and so I’ve done my best to
reassure them.  In a few days they will see that they can get by on much less water than they think
they need -- yet another way to experience for themselves the constraints space exploration will
place on early pioneers.
 
-----------------------------

CommanderCheckIn Report
01/09/2014
Paula Crock
Crew Physical Health:
Crew member with strained back is recovering and still performing
duties as assigned.  All other crew are well.

Eva Depart Time:
N/A
Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:
Water Delivery
Rep Sched:
Commander Check-In
Engineering Report
Science Report (GreenHab)
Future EVA Plan
Daily Photos
Journalist Report
Commander Report
Tomorrow Plan:
We have an EVA planned for tomorrow as well as some EVA suit cleaning.
The radio telescope may be functional starting tomorrow.
Inventory:
Added 4 4-roll packages of TP and 2 boxes of 22 count plastic trash bags
Support Req:
N/A
Miscelleneous:
D.G. said he will be out of town until Tuesday.